OpenStudy (anonymous):

Can someone tell me how to find the real and complex roots of this polynomial?

OpenStudy (anonymous):

What polynomial?

OpenStudy (anonymous):

This one: \[x^4 - 41x^2 = -400\]

OpenStudy (anonymous):

Sorry, my internet is really slow.

OpenStudy (anonymous):

Any ideas?

OpenStudy (dtan5457):

You can graph it and find the zeros.

OpenStudy (dtan5457):

x^4-41x^2+400=0 See where it touches the x intercepts?

OpenStudy (anonymous):

But is there any way I can solve it through factoring? I have to show my work for it.

OpenStudy (anonymous):

Could I maybe split the 41x^2 into two numbers?

Directrix (directrix):

@RadioEdit Picking up from above, you asked if x^4-41x^2+400=0 will factor (over the set of Integers, I assume). Yes, it will. Look for two numbers that multiply to 400 and add to -41. Hint: When you factor x^4 - 41x^2 + 400 = 0, you should get two binomial factors, each of which is the difference of two squares.